i am supposed to take a test for college that has linear equations such as c+(4-3c)-2=0 that was on a practice test they gave to me i do not remember how to solve this and i have looked up how to solve it and cannot figure it out based on what i found please just explain how to solve problems like this so i can do my best at the test thank you very much
Answer by Theo(13342)

equation is:
c + (4-3c) - 2 = 0
remove parentheses to get:
c + 4 - 3c - 2 = 0
combine like terms to get:
-2c + 2 = 0
subtract 2 from both sides of the equation to get:
-2c = -2
divide both sides of the equation by -2 to get:
c = -2/-2 = 1
answer should be c = 1
replace c with 1 in your original equation to get:
c + (4-3c) - 2 = 0 becomes:
1 + (4-3*1) - 2 = 0 which becomes:
1 + (4-3) - 2 = 0 which becomes:
1 + 1 - 2 = 0 which becomes:
2 - 2 = 0 which becomes:
0 = 0 confirming that the value of c = 1 is good.